# Going Proactive and Explanatory Against Malware Concept Drift

### v1

@@ 6.2 Drift Detection
Table 6: Explanation evaluation results on the Drebin dataset with CBP and DRR metrics. The columns labeled f0 to f7 represent the held-out families, ordered by their sample size in descending order.
@@
Table 7: Evaluation of feature-level explanations.
6.3 Drift Adaptation
@@
Evaluation Results. Table 4 shows the drift adaptation results across different labeling budgets on all datasets and classifiers. Our method outperforms the baseline in different settings without exception. Firstly, regarding performance improvements with different classifiers, we observed significant enhancements. On the Drebin dataset, the improvements are 82.0%, 69.3%, and 19.9% for Drebin, Mamadroid, and Damd classifiers, respectively. On the Malradar dataset, the improvements are even more pronounced at 171.9%, 15.9%, and 42.1% for the same classifiers.
Secondly, when considering varying human labeling budgets (10, 20, 30, 40, 100), the improvements on the Drebin dataset are 238.1%, 52.8%, 21.1%, 15.8%, and 6.5%, respectively. Similarly, on the Malradar dataset, the enhancements are 98.7%, 94.2%, 23.3%, 23.6%, and 5.5% for these respective budgets. A key finding is that the smaller the budget, the greater the improvement. This suggests that in scenarios prioritizing human analysis, our method can significantly reduces labeling budgets. For example, in order to achieve an accuracy score of 0.9 on the Drebin feature of the Drebin dataset, our method only needs 20 budgets, while the baseline requires 100, saving 80% of the analysis cost. On the Drebin feature of the Malradar dataset, the baseline method cannot achieve an accuracy score of 0.9 even if it uses 100 budgets, while our method can achieve an accuracy of 0.913 with the budget 20.
When considering the presence of labeling noise, the pivotal finding from Table 5 is that our method consistently achieves stable and higher F1-scores and accuracy scores compared to the baseline. Moreover, when the noise level increases, such as to 6%, the accuracy score improvement can reach 28.4%. This could be attributed to the inclusion of an explanation revision step during model updates, making our method more robust and resilient to noise.
6.4 Drift Explanation
Baseline and Metric. We consider three baseline methods for drift explanation, adapted to generate both concept-level and traditional feature-level explanations: 1) a random baseline that randomly selects features or concepts as important; 2) a gradient-based explainer that utilizes Integrated Gradients (IG) [63], adapted here to analyze drifting scores derived from our detector. Originally designed to attribute a deep network’s predictions to its input features, IG has demonstrated effectiveness in explicating supervised security applications [70, 27]. In this context, it is tailored to focus on the drifting scores derieved on model predictions, and the reference sample $\mathbf{x}_{r}$ serves as the baselines in its method; 3) the state-of-the-art CADE explainer used with the CADE detector, as described in Section 5.3. Since this method is not inherently designed for concept-space explanations, we adapt it to our detector for generating explanations in the concept space.
@@ 6.4 Drift Explanation
Evaluation Results. Table 6 illustrates the evaluation results of concept-based explanations generated by different methods. Comparing to three baseline explainers, i.e., Random, Dri-IG and $\text{CADE}^{\text{+}}$, Dream surpasses them by 172.1%, 97.2% and 11.5% on the CBP metric, and by 254.2%, 143.3% and 1.4% on the DRR metric. Notably, $\text{CADE}^{\text{+}}$ is included in this comparison as it utilizes our detector for concept-based explanations, extending beyond its original method’s capabilities. To examine the efficacy in areas typically addressed by common methods, Table 7 presents the outcomes for feature-level explanation evaluations. Here, our method, despite not being primarily designed for feature-level explanations, outperforms all baselines, including the state-of-the-art CADE explainer, benefiting from our method’s sensitive capture of deviations. Additionally, an intriguing observation is the contrast in CADE’s performance: its CBP value is 24.1% lower than that of Dri-IG, yet its DRR value is 61.9% higher. This discrepancy highlights the greater reliability of CBP over DRR in evaluating drift explainations. Combining the results from both tables, our explainer yields more substantial improvements in concept space than in feature space. For instance, while our method achieves a 45.2% increase in mean CBP over Dri-IG at the feature level, it shows a remarkable 97.2% increase at the concept level, highlighting the effectiveness of our design in concept space.
7 Extended Evaluation
@@
Evaluation Method. In accordance with the evaluation practices in Section 4.1, we also extend the baselines by enhancing the HCC method with two specific adaptations: 1) the removal of the surrogate classifier $\mathbf{M}_{s}$ and its integration into the original classifier using the autoencoder structure; 2) the augmentation of the previous cross-entropy based pseudo loss item with our newly proposed NCE method, as in Equation 11. Theses modifications firstly render the HCC method fully model sensitive for detecting drift of external classifiers, and secondly make the uncertainty measurement more effective based on our observations in Section 5.2.
Evaluation Results. As presented in Table 8, our modification achieves significant enhancements in the HCC method’s detection performance. This is evident both in terms of the integrated method and its two individual pseudo-loss detectors. Notably, there is an average performance increase from $9.43\%$ to $10.62\%$. The contrastive-based pseudo-loss detector, previously the least effective, now surpasses all prior baselines with the most substantial improvement. This enhancement can be primarily attributed to the model-sensitive training approach. More importantly, our method demonstrates superior performance overall. Compared to existing methods, we achieve an $18.57\%$ improvement. Even when comparing against the improved HCC method, our approach maintains an $8.21\%$ lead. This superior performance is largely due to our unique approach in concept-based training and the data-autonomous design of our drift detection model.
@@
Table 8: The enhancement of intra-class drift detection with our detection framework.
8 Discussion
@@
Concepts and Annotators. We define malicious behaviors as concepts that effectively handles inter-class drift. To achieve comprehensive drift adaptation, we can explore border concepts. For example, the benign functionality can be included to better deal with intra-class drift. Furthermore, in the active learning framework, the human annotation step can be done in a more intelligent manner. Future efforts may utilize dynamic analysis tools [18, 1, 42] to automatically output behaviors or employ LLM models [10, 6] to generate higher-level contents.
9 Related Work
In this section, we discuss four additional areas of related work to complement Section 2.
@@ 10 Conclusion
Appendix A Clarity of Notation
The notation in this paper is meticulously chosen and maintains consistency in the exposition of mathematical frameworks for malware concept drift handling. To ensure clarity, we adopt the following key formatting conventions:
  • Sets are denoted with calligraphic script (e.g., $\mathcal{X},\mathcal{Y}$).
  • Vectors and matrices are denoted by boldface letters to signify their multi-dimensional nature (e.g., $\mathbf{x}$).
  • Scalars and elements of sets are represented in standard italicized font (e.g., $y,m,n$).
  • Constant values, such as the number of malware families and the number of layers in a model, are represented as uppercase italicized characters (e.g., $C,K$).
  • High-level functions, such as the predictive model $\mathbf{M}$, along with the Drift Detector $\mathbf{D}$, the Drift Explainer $\mathbf{E}$, and the Drift Adaptor $\mathbf{A}$, are denoted by boldface letters, emphasizing their functional significance.
  • Lowercase three-letter abbreviations are reserved for unit functions within the primary processes (e.g., $unc$ for uncertainty estimation, $ncm$ for nonconformity scoring).
To facilitate reading, we list these symbols, as well as their descriptions and mappings in Table 9.
Symbol Description and Mapping
$\mathcal{X},\mathbf{x}$ Feature space and an input instance
$\mathcal{X}\subseteq\mathbb{R}^{p\times q},\mathbf{x}\in\mathcal{X}$
$\mathcal{Y},y$ Label space and the true label for $\mathbf{x}$
$\mathcal{Y}:=\{0,1\}\text{ or }\{1,2,\ldots,C\},y\in\mathcal{Y}$
$\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}$ Training and test datasets
$\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}\subseteq\mathcal{X}\times\mathcal{Y}$
$\mathbf{M}$ The target classifier
$\mathbf{M}:\mathcal{X}\rightarrow\mathcal{P}(\mathcal{Y})$
$unc,ncm$ Uncertainty estimation and nonconformity scoring functions
$unc:\mathcal{X}\times\mathbf{M}\rightarrow\mathbb{R}$, $ncm:\mathcal{X}\times\mathcal{Y}\times\mathcal{D}_{\text{train}}\rightarrow\mathbb{R}$
$\mathbf{D}$ Drift scoring function
$\mathbf{D}:\mathcal{X}\times\mathcal{Y}\times\mathbf{M}\times\mathcal{D}_{\text{train}}\rightarrow\mathbb{R}$
$\mathbf{A}$ Model updating function
$\mathbf{A}:\mathbf{M}\times\mathcal{D}_{\text{test}}\times\mathcal{E}\rightarrow\mathbf{M}^{\prime}$
Table 9: Notation table
Appendix B Malware Dataset
Family # sample (#) behavior Time
[…]
Table 10: Drebin (top) and Malradar (bottom) datasets. In the third column, the number of existed behaviors is in parentheses, followed by their indices, with the specific meaning provided in Appendix B.
As shown in Table 10, we use two datasets for malware family classification. Specifically, we select 8 families for each of the Drebin [75] and the MalRadar [69] dataset, where the family selection adhers to the same criteria used in the CADE paper. The Drebin dataset spans the years from 2010 to 2012, offering insights into malware characteristics from ealy period. To capture more recent trends in the evolving malware landscape, the Malradar dataset [69] covers the period from 2015 to 2021.
We use the malicious behaviors defined in Malradar paper, where $\text{b}_{0}$-$\text{b}_{9}$ correspond to privacy information stealing ($\text{b}_{0}$), SMS/phone calls ($\text{b}_{1}$), remote control ($\text{b}_{2}$), bank stealing ($\text{b}_{3}$), ransom ($\text{b}_{4}$), abusing accessibility ($\text{b}_{5}$), privilege escalation ($\text{b}_{6}$), stealthy download ($\text{b}_{7}$), ads ($\text{b}_{8}$), and premium service ($\text{b}_{9}$). To ensure consistency in behavioral analysis, we augment the Drebin dataset, which originally lacks behavior labels, with these same behaviors. This extension involves extrapolating the labels based on Malradar’s definitions, supplemented by expert analysis and validation through GPT [48]. Note that the concept labels were assigned with minimal additional effort, and it is uniform within each family. In cases where certain behaviors in the Malradar dataset are specific to particular variants, we accommodate this by setting the relevant valid label mask $m_{e}$ to zero. Despite the simplicity, this approach is in line with an active learning setting that typically utilizes limited human effort.
For the Drebin dataset, the time-split is performed across all remaining families, as detailed in the CADE paper. For the Malradar dataset, its extensive time range and emergence of new families in recent years necessitate a tailored splitting approach. The previous split would classify newer families, such as EventBot which emerged in 2020, entirely into the testing set, potentially introducing unintended drift. To avoid this, we adjust the strategy by performing the split on a per-family basis within this dataset.
Appendix C High-level Human Feedback
The high-level abstraction of concept-based explanations allows feedback to be more generalized, potentially encompassing instances of a specific class and reducing the effort required for annotating explanations. An example of such generalized semantic feedback is: {mdframed}[style=mystyle]
$\begin{split}\forall mal\,\,\,\,&\texttt{Family}(mal,\texttt{GhostCtrl})\Rightarrow\\ &\texttt{Behaviors}(mal)\Rightarrow(\texttt{PrivacyStealing}(mal)\\ &\land\texttt{SMSCALL}(mal)\land\texttt{RemoteControl}(mal)),\end{split}$
which implies that for all drifting samples in the GhostCtrl family, the expected behaviors include PrivacyStealing, SMSCALL, and RemoteControl, acknowledging that only some variants may exhibit the Ransom behavior.
Appendix D Autoencoder Architecture
For the Drebin feature, the autoencoder used is a form of Tabular Autoencoder consisting of two dense layers in both the encoder and decoder components. The architecture features a hidden dimension size of $512$ and an encoding dimension of $32$. In the case of the Mamadroid feature, the autoencoder is similarly structured as a Tabular Autoencoder, but with a larger hidden dimension of $2048$ and an encoding dimension of $128$. This expanded architecture accommodates the more complex nature of the Mamadroid feature set. For the Damd feature, the autoencoder is a specialized convolutional Text Autoencoder. This autoencoder works in conjunction with an embedding layer to facilitate reconstruction on numerical data. The encoder incorporates a convolutional layer with a kernel size of $3$ and $64$ filters, followed by a global max pooling layer. The decoder comprises a dense layer and a convolutional transpose layer that transform the data back into the embedding dimension, maintaining the same kernel size. This configuration is tailored to effectively handle the sequential nature of this feature set.
Appendix E Concept-based Explanation
The drift explainer in our system can explain OOD data, identifying concepts which contributes most to the drifting. As in Figure 4, we make an illustration of the drift explanations on the Drebin dataset.
Figure 4: The concept-based drift explanation heatmap on Drebin dataset, where the horizontal axis represents family names, and the vertical axis represents the proportion of explanation values that are positive for each behavior.
Hold-out Baseline Ours Imp.
FakeInstaller 0.472 0.999 111.8%
DroidKungFu 0.496 1.000 101.8%
Plankton 0.516 1.000 93.9%
GingerMaster 0.526 1.000 90.2%
BaseBridge 0.527 1.000 89.8%
Iconosys 0.532 1.000 87.9%
Kmin 0.532 1.000 87.8%
FakeDoc 0.532 1.000 87.8%
Table 11: Concept classification accuracy for ID test data.
Besides drift explanation, our approach leverages an autoencoder intermediary to generate concept-based explanations, providing a novel tool for in-depth analysis of in-distribution data. The challenge in offering high-level explanations, beyond mere feature attribution, persists even in non-drifting scenarios. Current methodologies tend to abstract explanations from features [54, 27], but they cannot readily translate into behavioral insights.
To establish a baseline for concept-based explanations, we frame it as a series of binary classification tasks, each corresponding to a different explicit concept. Specifically, we adapt the output layer of the classifier to yield sigmoid probabilities and conduct fine-tuning over 100 retraining epochs. This ensures that each classifier achieves an accuracy of over 0.99 during training. We then select in-distribution data from the test dataset—specifically, data that corresponds to classes previously encountered during training—and assess the concept classification accuracy across all concepts.
We conduct the experiments with the Drebin dataset, and as detailed in Table 11, the results reveal that our method achieves near-perfect accuracy on the non-drifting test dataset, approaching $100\%$. This represents a significant $93.9\%$ improvement over the baseline across all models, indicating strong stability on in-distribution data. However, we observed a stark decrease in concept accuracy on out-of-distribution data, with average accuracies around $66\%$ for both the baseline and our method. Interestingly, the baseline shows a slight improvement in this context, possibly due to its conservative nature, leading to more frequent negative outputs. These ou emphasize the necessity of updating both the detector and classifier during drift adaptation, a key aspect of our drift adaptor’s design.
(a) Transcendent, Drebin
(b) Transcendent, Mamadroid
(c) Transcendent, Damd
(d) CADE, Drebin
(e) CADE, Mamadroid
(f) CADE, Damd
(g) Probability, Drebin
(h) Probability, Mamadroid
(i) Probability, Damd
(j) (Ours) DREAM, Drebin
(k) (Ours) DREAM, Mamadroid
(l) (Ours) DREAM, Damd

### v2

@@ VI-B Drift Detection
TABLE VI: Explanation evaluation results on the Drebin dataset with CBP and DRR metrics. The columns labeled f0 to f7 represent the held-out families, ordered by their sample size in descending order.
@@
TABLE VII: Evaluation of feature-level explanations.
VI-C Drift Adaptation
@@
Evaluation Results. Table IV shows the drift adaptation results across different labeling budgets on all datasets and classifiers. Our method outperforms the baseline in different settings without exception. Firstly, regarding performance improvements with different classifiers, we observed significant enhancements. On the Drebin dataset, the improvements of F1-score are 95.2%, 82.7%, and 22.6% for Drebin, Mamadroid, and Damd classifiers, respectively. On the Malradar dataset, the improvements are 96.3%, 11.6%, and 39.2% for the same classifiers.
Secondly, when considering varying human labeling budgets (10, 20, 30, 40, 100), the improvements of F1-score on the Drebin dataset are 238.1%, 52.8%, 21.1%, 15.8%, and 6.5%, respectively. Similarly, on the Malradar dataset, the enhancements are 98.7%, 94.2%, 23.3%, 23.6%, and 5.5% for these respective budgets. A key finding is that the smaller the budget, the greater the improvement. This suggests that in scenarios prioritizing human analysis, our method can significantly reduces labeling budgets. For example, in order to achieve an accuracy score of 0.9 on the Drebin feature of the Drebin dataset, our method only needs 20 budgets, while the baseline requires 100, saving 80% of the analysis cost. On the Drebin feature of the Malradar dataset, the baseline method cannot achieve an accuracy score of 0.9 even if it uses 100 budgets, while our method can achieve an accuracy of 0.913 with the budget 20.
When considering the presence of labeling noise, the pivotal finding from Table V is that our method consistently achieves stable and higher F1-scores and accuracy scores compared to the baseline. Moreover, when the noise level increases, such as to 6%, the accuracy score improvement can reach 28.4%. This could be attributed to the inclusion of an explanation revision step during model updates, making our method more robust and resilient to noise.
VI-D Drift Explanation
Baseline and Metric. We consider three baseline methods for drift explanation, adapted to generate both concept-level and traditional feature-level explanations: 1) a random baseline that randomly selects features or concepts as important; 2) a gradient-based explainer that utilizes Integrated Gradients (IG) [52], adapted here to analyze drifting scores derived from our detector. Originally designed to attribute a deep network’s predictions to its input features, IG has demonstrated effectiveness in explicating supervised security applications [53, 18]. In this context, it is tailored to focus on the drifting scores derieved on model predictions, and the reference sample $\mathbf{x}_{r}$ serves as the baselines in its method; 3) the state-of-the-art CADE explainer used with the CADE detector, as described in Figure 4. Since this method is not inherently designed for concept-space explanations, we adapt it to our detector for generating explanations in the concept space.
@@ VI-D Drift Explanation
Evaluation Results. Table VI illustrates the evaluation results of concept-based explanations generated by different methods. Comparing to three baseline explainers, i.e., Random, Dri-IG and $\text{CADE}^{\text{+}}$, Dream surpasses them by 172.1%, 97.2% and 11.5% on the CBP metric, and by 254.2%, 143.3% and 1.4% on the DRR metric. Notably, $\text{CADE}^{\text{+}}$ is included in this comparison as it utilizes our detector for concept-based explanations, extending beyond its original method’s capabilities. To examine the efficacy in areas typically addressed by common methods, Table VII presents the outcomes for feature-level explanation evaluations. Here, our method, despite not being primarily designed for feature-level explanations, outperforms all baselines, including the state-of-the-art CADE explainer, benefiting from our method’s sensitive capture of deviations. Additionally, an intriguing observation is the contrast in CADE’s performance: its CBP value is 24.1% lower than that of Dri-IG, yet its DRR value is 61.9% higher. This discrepancy highlights the greater reliability of CBP over DRR in evaluating drift explainations. Combining the results from both tables, our explainer yields more substantial improvements in concept space than in feature space. For instance, while our method achieves a 45.2% increase in mean CBP over Dri-IG at the feature level, it shows a remarkable 97.2% increase at the concept level, highlighting the effectiveness of our design in concept space.
VII Extended Evaluation
@@
Evaluation Method. In accordance with the evaluation practices in Section IV-B, we also extend the baselines by enhancing the HCC method with two specific adaptations: 1) the removal of the surrogate classifier $\mathbf{M}_{s}$ and its integration into the original classifier using the autoencoder structure; 2) the augmentation of the previous cross-entropy based pseudo loss item with our newly proposed NCE method, as in Equation 11. Theses modifications firstly render the HCC method fully model sensitive for detecting drift of external classifiers, and secondly make the uncertainty measurement more effective based on our observations in Section V-B.
Evaluation Results. As presented in Table VIII, our modification achieves significant enhancements in the HCC method’s detection performance. This is evident both in terms of the integrated method and its two individual pseudo-loss detectors. Notably, there is an average performance increase from $9.43\%$ to $10.62\%$. The contrastive-based pseudo-loss detector, previously the least effective, now surpasses all prior baselines with the most substantial improvement. This enhancement can be primarily attributed to the model-sensitive training approach. More importantly, our method demonstrates superior performance overall. Compared to existing methods, we achieve an $18.57\%$ improvement. Even when comparing against the improved HCC method, our approach maintains an $8.21\%$ lead. This superior performance is largely due to our unique approach in concept-based training and the data-autonomous design of our drift detection model.
@@
TABLE VIII: The enhancement of intra-class drift detection with our detection framework.
VIII Discussion
@@
Concepts and Annotators. We define malicious behaviors of Android malware as concepts that effectively handles inter-class drift. To achieve comprehensive drift adaptation, we can explore border concepts. For example, including concepts of benign functionality helps address intra-class drift, while adding specific behaviors accommodates other platforms [48]. Furthermore, in the active learning framework, the human annotation step can be done in a more intelligent manner. Future efforts may utilize dynamic analysis tools [59, 60, 61] to automatically output behaviors or employ LLM models [62, 63] to generate higher-level contents.
IX Related Work
In this section, we discuss four additional areas of related work to complement Section II.